OpenStudy (anonymous):

what is the factored form of 2x squared plus 20x plus 50

jhonyy9 (jhonyy9):

like a first step you can factorizing out 2 and will get 2(x^2 +10x +25) = x^2 +10x +25 = x*x +5x +5x + 5*5 = x(x+5) +5(x+5) = (x+5)(x+5) = (x +5)^2 so than 2x^2 +20x +50 = 2(x+5)^2 hope so much that is understandably good luck